A city park needs a new fence. The fence costs $10.00 per meter, and the park is 68 meters by 87 meters. How much will it cost to put up a new fence around the park?
Answer by macston(5194)

.
The length of the fence will be the perimeter of the park.
.
L=length of park=87m; W=width of park=68m; P=perimeter of park
.
P=2(L+W)
P=2(87m+68m)
P=2(155m)
P=310m
.
The length of the fence will be 310 meters.
.
Cost=($10.00)(perimeter)=($10.00)(310m)=$3100
.
ANSWER: the fence will cost $3100.
.